# Robust Loss Functions for Training Decision Trees with Noisy Labels

### 20 Dec 2023 v1

@@ 1 Introduction
The remainder of this paper is organized as follows. Section 2 provides a more detailed discussion on related work and Section 3 presents some preliminary concepts. Section 4 presents the conservative losses and their robustness properties. Section 5 describes our distribution-based robust loss framework and the negative exponential loss. Section 6 presents details on experimental settings and results. Our source code is available at https://github.com/jonathanwilton/RobustDecisionTrees.
2 Related Work
@@ 3 Preliminaries
In the recursive greedy risk minimization perspective, a node is split in a greedy way to minimize the empirical risk $\widehat{R}({\bm{g}};{\cal{D}})$. If ${\bm{g}}$ predicts a constant $\widehat{{\bm{y}}}$ on a subset ${\cal{S}}\subseteq{\cal{D}}$ of the training examples, then the contribution to the empirical risk is the partial empirical risk $\widehat{R}(\widehat{{\bm{y}}};{\cal{S}}):=\sum_{({\bm{x}},{\bm{y}})\in{\cal{S}}}\ell(\widehat{{\bm{y}}},{\bm{y}})/|{\cal{D}}|$. Denote by $\widehat{{\bm{y}}}^{*}:=\mathrm{argmin}_{\widehat{{\bm{y}}}\in\Delta}\widehat{R}(\widehat{{\bm{y}}};{\cal{S}})$ the optimal constant probability vector prediction, with minimum partial empirical risk $\widehat{R}^{*}({\cal{S}}):=\widehat{R}(\widehat{{\bm{y}}}^{*};{\cal{S}})$. The minimum partial empirical risk can be interpreted as an impurity measure. In fact, it is the Gini impurity and the entropy impurity (up to a multiplicative constant) when the loss is the MSE loss and the CE loss, respectively.
@@
4 Conservative Losses
We first examine the impurities corresponding to various loss functions, including both standard loss functions and robust ones, in Theorem 1. Parts $(a)$, $(b)$ and $(c)$ are shown in previous works (Breiman et al. 1984; Painsky and Wornell 2018; Yang, Gao, and Li 2019; Wilton et al. 2022) but included for completeness. All proofs are given in the appendices. Unless otherwise stated, we shall use ${\cal{D}}$ to denote an arbitrary (possibly noisy) set of input-output pairs, and ${\cal{S}}$ a subset of ${\cal{D}}$, in this section.
@@ 4 Conservative Losses
5 Distribution Losses
We introduce a new approach for constructing robust loss functions that can adapt to the noise level. This allows us to address a limitation of the conservative losses as pointed out in Section 4: while conservative losses are robust against label noise, they can lead to underfitting in low noise situations.
@@ 5 Distribution Losses
Figure 1: Left: NE loss as a function of the margin $y\widehat{y}$. Right: Cross entropy (also GCE impurity with $q=0$), GCE impurities for $q=0.2,0.5,0.8$, the misclassification impurity (also GCE impurity with $q=1$ and NE impurity with $\lambda=1$), and NE impurities for $\lambda=0.2,0.5,0.8$. The impurities have been scaled for better comparison.
@@
Note that in the definition of the NE impurity above, $\widehat{y}$ ranges over ${\mathbb{R}}$ instead of over the 1-simplex. This is because we have chosen the prediction $\widehat{y}$ to be a positiveness score, instead of a two-dimensional vector representing a probability distribution, as done in Section 3.
@@ 6 Experiments
We consider some commonly used datasets from UCI including Covertype (Blackard 1998), 20News (Lang 1995), Mushrooms (mis 1987), as well as the MNIST digits (LeCun et al. 1998), CIFAR-10 (Krizhevsky 2009) and UNSW-NB15 (Moustafa and Slay 2015). Datasets were selected due to diversity in number of samples, number of features, number of classes, type (image, tabular, text), and domain (computer vision, cyber security, cartography). We look at both multiclass ($K>2$) and binary ($K=2$) classification problems for each dataset. Mushrooms datasets and ranking loss excluded from multiclass classification experiments due to Mushrooms being a binary classification dataset, and ranking loss only defined for binary classification problems. Table 1 is a summary of the benchmark datasets.
@@
We trained each model on 5 random noisy training sets to account for randomness in training labels and learning algorithms. We report the average test set accuracies and 2x standard deviation bands. Additional results can be found in Appendices L and K.
@@ Appendix I Proof of Theorem 8
We performed additional experiments on CIFAR-10 using pre-trained image embeddings from VGG11111https://github.com/huyvnphan/PyTorch˙CIFAR10. The input to the last classifier module in the network was taken to be the new feature vectors. Other processing steps are the same as in the other experiments.

### 23 Jan 2024 v2

@@ Introduction
The remainder of this paper is organized as follows. We first provide a more detailed discussion on related work, followed by some preliminary concepts. We then present the conservative losses and their robustness properties. After that, we describe our distribution-based robust loss framework and the negative exponential loss. Finally, we present details on experimental settings and results, with a brief conclusion. Our source code is available at https://github.com/jonathanwilton/RobustDecisionTrees.
Related Work
@@
In the recursive greedy risk minimization perspective, a node is split in a greedy way to minimize the empirical risk $\widehat{R}({\bm{g}};{\cal{D}})$. If ${\bm{g}}$ predicts a constant $\widehat{{\bm{y}}}$ on a subset ${\cal{S}}\subseteq{\cal{D}}$ of the training examples, then the contribution to the empirical risk is the partial empirical risk $\widehat{R}(\widehat{{\bm{y}}};{\cal{S}}):=\sum_{({\bm{x}},{\bm{y}})\in{\cal{S}}}\ell(\widehat{{\bm{y}}},{\bm{y}})/|{\cal{D}}|.$ Denote by $\widehat{{\bm{y}}}^{*}:=\mathrm{argmin}_{\widehat{{\bm{y}}}\in\Delta}\widehat{R}(\widehat{{\bm{y}}};{\cal{S}})$ the optimal constant probability vector prediction, with minimum partial empirical risk
$\widehat{R}^{*}({\cal{S}}):=\widehat{R}(\widehat{{\bm{y}}}^{*};{\cal{S}}).$
The minimum partial empirical risk can be interpreted as an impurity measure. In fact, it is the Gini impurity and the entropy impurity (up to a multiplicative constant) when the loss is the MSE loss and the CE loss, respectively.
@@
Conservative Losses
We first examine the impurities corresponding to various loss functions, including both standard loss functions and robust ones, in Theorem 1. Parts $(a)$, $(b)$ and $(c)$ are shown in previous works (Breiman et al. 1984; Painsky and Wornell 2018; Yang, Gao, and Li 2019; Wilton et al. 2022) but included for completeness. All proofs are given in the appendices.  [^1] The appendices are available in the full version of the paper at https://github.com/jonathanwilton/RobustDecisionTrees. Unless otherwise stated, we shall use ${\cal{D}}$ to denote an arbitrary (possibly noisy) set of input-output pairs, and ${\cal{S}}$ a subset of ${\cal{D}}$, in this section.
@@
Distribution Losses
We introduce a new approach for constructing robust loss functions that can adapt to the noise level. This allows us to address a limitation of the conservative losses as pointed out in the previous section: while conservative losses are robust against label noise, they can lead to underfitting in low noise situations.
@@
Figure 1: Left: NE loss as a function of the margin $y\widehat{y}$. Right: Blue lines are NE impurities with their $\lambda$ values provided in the legend. The special case with $\lambda=1$ is denoted by CL (conservative loss). Red lines are GCE impurities with their $q$ values provided in the legend. The special case with $q=0$ is denoted by CE (cross entropy). The impurities have been scaled for better comparison.
@@
Note that in the definition of the NE impurity above, $\widehat{y}$ ranges over ${\mathbb{R}}$ instead of over the 1-simplex. This is because we have chosen the prediction $\widehat{y}$ to be a positiveness score, instead of a two-dimensional vector representing a probability distribution, as done in the preliminaries section.
@@
We consider some commonly used datasets from UCI including Covertype (Blackard 1998), 20News (Lang 1995), Mushrooms (Audubon Society Field Guide 1987), as well as the MNIST digits (LeCun et al. 1998), CIFAR-10 (Krizhevsky 2009) and UNSW-NB15 (Moustafa and Slay 2015). Datasets were selected due to diversity in number of samples, number of features, number of classes, type (image, tabular, text), and domain (computer vision, cyber security, cartography). We look at both multiclass ($K>2$) and binary ($K=2$) classification problems for each dataset. Mushrooms datasets and ranking loss excluded from multiclass classification experiments due to Mushrooms being a binary classification dataset, and ranking loss only defined for binary classification problems. Table 1 is a summary of the benchmark datasets.
@@
We trained each model on 5 random noisy training sets to account for randomness in training labels and learning algorithms. We report the average test set accuracies and 2x standard deviation bands. Additional results can be found in Appendices K and L.
@@ Appendix I Proof of Theorem 8
We performed additional experiments on CIFAR-10 using pre-trained image embeddings from VGG11222https://github.com/huyvnphan/PyTorch˙CIFAR10. The input to the last classifier module in the network was taken to be the new feature vectors. Other processing steps are the same as in the other experiments.